# Federated Learning Clients Clustering with Adaptation to Data Drifts

### v1

@@ 6.1. End-to-end Training Performance
Figure 6. Fielding with different client selection strategies on FMoW dataset.
6.2. Compatibility with Other Optimizations
To demonstrate that Fielding is agnostic to client selection and model aggregation strategies, we evaluate Fielding’s performance on FMoW together with various complementary optimizations. In Figure 6 and Figure 7, we use the algorithm name alone to denote the baseline where we train one global model and use “+ Fielding ” to denote running the algorithm atop Fielding.
We use the Oort selection algorithm (Lai et al., 2021) and a distance-based algorithm prioritizing clients closer to the distribution center as client selection examples. As shown in Figure 6, Fielding improves the final average test accuracy by 6.0% and 4.5% while reaching the target accuracy 2.62×2.62$\times$2.62 × and 1.17×1.17$\times$1.17 × faster. Note that Oort selection actively incorporates clients otherwise filtered out due to long response time, leading to a longer average round time.
We also show that Fielding works well with aggregation algorithm FedYogi (Reddi et al., 2021) and q-FedAvg (Li et al., 2020c) in Figure 7. Fielding improves the final accuracy by 5.9% and 4.9% while giving a 1.34×1.34$\times$1.34 × and 1.25×1.25$\times$1.25 × speedup respectively.
6.3. Robustness and Sensitivity Analysis
@@
Figure 7. Fielding with different FL algorithms on FMoW dataset.
Figure 8. Fielding is robust to malicious clients.
Figure 9. Fielding offers gains across heterogeneity degrees.
Figure 10. Fielding offers gains across heterogeneity degrees.
@@
7. Related Work and Discussion
Clustered FL Previous clustering solutions (Ghosh et al., 2020; Duan et al., 2021; Liu et al., 2023; Zhang et al., 2024; Jothimurugesan et al., 2023) have achieved success at bounding intra-cluster heterogeneity and improving model performance. However, their improvements are compromised by real-world data drifts as they either assume stable environments or handle drifts insufficiently.
@@ 7. Related Work and Discussion
Heterogeneity-aware FL. Recent works have tackled heterogeneity challenges through client selection, workload scheduling, and update corrections. Oort considers both statistical and system utility and designs an exploration-exploitation strategy for client selection (Lai et al., 2021). PyramidFL adjusts the number of local iterations and parameter dropouts to optimize participants’ data and system efficiency (Li et al., 2022b). DSS-Edge-FL dynamically determines the size of local data used for training and selects representative samples, optimizing resource utilization while considering data heterogeneity (Serhani et al., 2023). MOON corrects local updates on clients by maximizing the similarity between representations learned by local models and the global model (Li et al., 2021).
@@
Drifts-aware FL. Adaptive-FedAvg is an FL algorithm that adapts client learning rate based on the model variability between consecutive rounds (Canonaco et al., 2021). CDA-FedAvg detects concept drifts and extends FedAvg for learning under such drifts. It defines a short-term and a long-term memory for each client and applies rehearsal using data in the lone-term memory when drifts happen (Casado et al., 2022). Master-FL proposes a multi-scale algorithmic framework that instructs clients to train over multiple time horizons with a suitable learning rate for each time chunk (Ganguly and Aggarwal, 2023).
@@
Fielding focuses on dynamic client clustering that is robust against potential data drifts. It is agnostic to other FL optimizations and should work seamlessly with the FL aggregation, selection, and scheduling algorithms above.
8. Conclusion
@@
We have presented Fielding, a novel clustering-based Federated Learning framework designed to handle data drifts with low overhead. Fielding addresses the challenges of data drift by re-clustering clients based on their label distribution vectors as their data distributions change, dynamically adjusting the number of clusters, and stabilizing responses to minor data shifts. Fielding is robust against malicious clients and varying degrees of heterogeneity, and it is compatible with various client selection and aggregation algorithms. In Fielding, clients register with a centralized coordinator that updates their metadata and assigns them to the nearest cluster based on their local data distribution. Before each training round, the coordinator determines how much each cluster center has shifted. If any cluster exceeds a predefined threshold, a global re-clustering process is initiated; otherwise, the updated cluster memberships are finalized. Experimental results demonstrate that Fielding improves the final test accuracy by 1.9% to 5.9% compared to existing methods and provides greater stability when facing real-world data drifts across different settings.
There are potential directions to further enhance Fielding. Our current approach requires labeled data, which may not be available in all contexts. Additionally, clients actively report their label distribution vectors to the coordinator, which introduces latent privacy concerns. In future work, we will focus on alternatives, such as inferring distribution vectors from gradients.
@@ Appendix A Proof of Convergence
1 input: the number of clients $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$
@@ 7             Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
8      Let $C_{1},\ldots,C_{K}$ be the $K$-center clustering of all clients based on representations from model $\tilde{\mathbf{x}}_{\tau}$
@@
10             $\tilde{\mathbf{c}}^{(k)}_{0}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$
@@ 17             $\mathbf{c}^{(k)}_{t+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{R}$
In this section, we analyze the performance of our framework in the case when the number of clusters $K$ is known and the number of local iterations is $1$. The analysis can be modified to handle local iterations and other variations based on the previous work, see e.g. Li et al. (2020a). In the algorithm, we recluster clients after every data drift event and use the $k$-center clustering instead of $k$-means clustering to simplify the proof.
We present the algorithm in Algorithm 4. After every data drift event, we first accept data changes for each client. After that, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster. After that, for several rounds, we perform standard Federated Averaging updates on each cluster: we sample $M$ clients, compute the stochastic gradient for each client, and update the cluster model with the sampled gradients.
@@
Let $f^{(i)}_{t}$ be the local function at each client $i$ at data drift event $t$. Then $f^{(i)}_{t}$ is $\mu$-strongly convex and $L$-smooth (has an $L$-Lipschitz gradient) for all $i$ and $t$.
The following are the standard assumptions on stochastic gradient.
@@
For each client $i$, at every data drift event $t$, we have access to stochastic gradient oracle $G^{(i)}_{t}$:
  • $\mathbb{E}\left[G^{(i)}_{t}(\mathbf{x})\right]=f^{(i)}_{t}(\mathbf{x})$ for all $\mathbf{x}$;
  • $\mathbb{E}\left\|G^{(i)}_{t}(\mathbf{x})-f^{(i)}_{t}(\mathbf{x})\right\|^{2}\leq\sigma^{2}$ for all $\mathbf{x}$.
The next assumption connects the objective value with client representations. Client representations can take various forms, and in our paper, we assume that the representation of the cluster is its label distribution.
@@
For a metric space $(X,\rho)$, let $r^{(i)}_{t}\in X$ be a representation of client $i$ at time $t$. Then there exists a constant $\theta$ such that $\left|f^{(i)}_{t}(\mathbf{x})-f^{(j)}_{t}(\mathbf{x})\right|\leq\theta\cdot\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)$ for all $i,j,t$.
@@
Finally, we assume that the clients are clusterable: there exist $K$ clusters so that within each cluster all points have similar representation.
@@
For completeness, we present the SGD convergence analysis for strongly convex functions. The following analysis is well-known, but, compared with the most standard analysis, it works directly with the loss functions instead of distances to the optimum, thus avoiding losing the condition number in the following bounds.
Lemma A.1.
@@
where in the last inequality we used $\eta\leq\frac{1}{L}$. Using the fact that $f$ is $\mu$-strongly-convex, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\geq 2\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))$, giving
@@
The above result shows how the objective improves with each round: namely, after $R$, the non-stochastic part of the loss improves by a factor which depends on $R$ exponentially. However, each data drift and each reclustering might potentially hurt the objective. Our next result bounds their effect on the loss.
Lemma A.2.
@@
$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})=(f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*}))+(f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*}))$
@@
(2) $f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})=f^{(i)}_{t+1}\Bigl{(}\operatorname*{avg}_{j\in\bar{C}_{\ell}}\mathbf{x}_{j}\Bigr{)}\leq\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j}),$
@@
where the last inequality holds since otherwise $f^{(i)}_{t}(\mathbf{c}_{k}^{*})>f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$, which is impossible since $\mathbf{c}_{k}^{*}$ is the minimizer of $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}$ and $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}(\mathbf{x}_{t}^{(i,*)})\leq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$ by Assumptions D and 3.
@@
Combining these results leads to our main convergence result.
Theorem A.3.
@@
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
@@
Next, by Lemma A.2, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering. Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
@@
$\leq\frac{1}{N}\sum_{k\in[K]}|C_{k}|\left(\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)+3\theta(\Delta+\delta)\right)(1-\eta\mu)^{R}+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
$=(1-\eta\mu)^{R}3\theta(\Delta+\delta)+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
@@
By Lemma A.1, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$. By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$. And finally, $f(\mathbf{x}_{0})-f(\mathbf{x}^{*})$ term is multiplied by $(1-\eta\mu)^{R}$ at every data drift event, giving factor of $(1-\eta\mu)^{TR}$ over $T$ data drift events. ∎
Appendix B Experiment settings

### v2

@@ 3.3 Robustness and sensitivity analysis
Figure 5: Fielding with different client selection strategies on FMoW dataset.
@@
Figure 6: Fielding with different FL algorithms on FMoW dataset.
Figure 7: Small-scale comparison against FedDrift on FMoW.
Figure 8: Fielding is robust to malicious clients.
Figure 9: Fielding offers gains across heterogeneity degrees.
Figure 10: Fielding improves model accuracy on static data.
@@
4 Related works
@@
Heterogeneity-aware FL. Recent works have also tackled heterogeneity challenges through client selection, workload scheduling, and update corrections. Oort considers both statistical and system utility and designs an exploration-exploitation strategy for client selection (Lai et al., 2021). PyramidFL adjusts the number of local iterations and parameter dropouts to optimize participants’ data and system efficiency (Li et al., 2022b). DSS-Edge-FL dynamically determines the size of local data used for training and selects representative samples, optimizing resource utilization while considering data heterogeneity (Serhani et al., 2023). MOON corrects local updates on clients by maximizing the similarity between representations learned by local models and the global model (Li et al., 2021).
@@
Drifts-aware FL. Adaptive-FedAvg handles concept drift by extending FedAvg with a learning rate scheduler that considers the variance of the aggregated model between two consecutive rounds (Canonaco et al., 2021). CDA-FedAvg detects concept drifts and extends FedAvg with a short-term and a long-term memory for each client. It applies rehearsal using data in the lone-term memory when drifts happen (Casado et al., 2022). Master-FL proposes a multi-scale algorithmic framework that trains clients across multiple time horizons with adaptive learning rates (Ganguly and Aggarwal, 2023). They are complementary to our work and can improve cluster models when minor drifts that don’t trigger global re-clustering happen.
@@
Fielding focuses on dynamic client clustering that is robust against potential data drift. It is agnostic to other FL optimizations and should work seamlessly with the FL aggregation, selection, and scheduling algorithms above.
Clustered FL. FlexCFL (Duan et al., 2021) and IFCA (Ghosh et al., 2020) adjust clusters by relocating shifted clients while maintaining a fixed number of clusters, operating under the assumption of low drift magnitude. Auxo (Liu et al., 2023) and FedAC (Zhang et al., 2024) re-cluster selected clients using gradient information or local models, delaying drift adjustments for unselected clients. FedDrift (Jothimurugesan et al., 2023) re-clusters all clients by broadcasting all cluster models every round to compute client local training loss, leading to significant costs. Appendix B discusses other prior CFL works without drift handling and the impact of drifts on static clusters.
5 Conclusion
@@
We presented Fielding, a clustering-based Federated Learning framework designed to handle multiple types of data drift with low overhead. Effective drift handling requires adapting to varying drift magnitudes while maintaining efficiency and supporting diverse drift types. To meet these goals, Fielding makes two key design choices: it combines per-client adjustment with selective global re-clustering and re-clusters all drifted clients using lightweight, drift-aware representations. Fielding dynamically adjusts the number of clusters and stabilizes the system’s response to minor drifts. It is robust to malicious clients and varying levels of heterogeneity, and remains compatible with a range of client selection and aggregation strategies. In Fielding, clients register with a centralized coordinator, which maintains metadata and assigns each client to the nearest cluster based on its local representation. Before each training round, the coordinator checks for shifts in cluster centers. If any center moves beyond a predefined threshold, global re-clustering is triggered; otherwise, the system proceeds with updated cluster assignments, significantly reducing the overall re-clustering cost and mitigating loss increase after re-clustering. Experimental results show that Fielding improves final test accuracy by 1.9%–5.9% over existing methods and provides greater stability under real-world data drift.
Limitations. There are potential directions to further enhance Fielding from the perspective of client representation. Reporting these representations – label distributions, embeddings, or gradients – poses potential privacy concerns. The differentially private approaches for private clustering are hard to apply in our settings since we perform clustering multiple times. Additionally, gradient representation, while being able to capture concept drifts, leads to significant overhead. We will address these concerns in future work.
@@ Appendix A Proof of convergence
1 input: the number of clusters $K$, the number of data drift events $T$, the number of sampled clients per round $M$, the number of rounds per data drift event $R$, reclustering threshold $\Delta$
@@ 7             Let $\mathbf{x}_{i}$ be the model corresponding to the cluster containing client $i$
8      Let $r^{(1)}_{t},r^{(2)}_{t},\ldots$ be the client representations
9       if there exists a cluster $C$ such that $\rho(r^{(i)}_{t},r^{(j)}_{t})>\Delta$ for clients $i,j\in C$ then
10             Let $C_{1},\ldots,C_{K}$ be the $K$-center clustering of all clients based on representations
11      else
12             Let $C_{1},\ldots,C_{K}$ be the current clusters
@@
14             $\tilde{\mathbf{c}}^{(k)}_{0}\leftarrow\operatorname*{avg}_{i\in C_{k}}\mathbf{x}_{i}$ // Cluster model is the average of clients’ models
@@ 21             $\mathbf{c}^{(k)}_{t+1}\leftarrow\tilde{\mathbf{c}}^{(k)}_{R}$
In this section, we analyze the performance of our framework in the case when the number of clusters $K$ is known and the number of local iterations is $1$. The analysis can be modified to handle local iterations and other variations based on the previous work, see e.g. Li et al. (2020a). To simplify the proof, in the algorithm, we make the following changes compared with the algorithm from the main body: 1) we use the $k$-center clustering instead of $k$-means clustering; 2) we perform reclustering when the intra-clsuter heterogeneity exceeds a certain threshold. We compare this threshold condition with the one from the main body in Section F.2.
We present the algorithm in Algorithm 4. After every data drift event, we first accept data changes for each client. After that, if the heterogeneity within some clsuter exceeds a certain threshold, we recluster the clients, and assign to each cluster a model by averaging models of all clients in the cluster. After that, for several rounds, we perform standard federated averaging updates on each cluster: we sample $M$ clients, compute the stochastic gradient for each client, and update the cluster model with the sampled gradients.
@@
Let $f^{(i)}_{t}$ be the local function at each client $i$ at data drift event $t$. Then, for all $i$ and $t$:
  • $f^{(i)}_{t}$ is $L$-smooth: $\|\nabla f^{(i)}_{t}(\mathbf{x})-\nabla f^{(i)}_{t}(\mathbf{x})\|\leq L\left\|\mathbf{x}-\mathbf{y}\right\|$ for all $\mathbf{x},\mathbf{y}$
  • $f^{(i)}_{t}$ satisfies $\mu$-Polyak-Łojasiewicz ($\mu$-PL) condition: $\|f^{(i)}_{t}(\mathbf{x})\|^{2}\geq 2\mu(f^{(i)}_{t}(\mathbf{x})-f^{(i)}_{t}(\mathbf{x}^{*}))$, where $\mathbf{x}^{*}$ is the minimizer of $f^{(i)}_{t}$.
The $\mu$-PL condition is a relaxation of the strong convexity assumption; both asumptions are standard in the federated learning literature (Cho et al., 7 29; Li et al., 2020a; Cho et al., 2020). The following are standard assumptions on stochastic gradient.
@@
For each client $i$, at every data drift event $t$, let $G^{(i)}_{t}$ be the stochastic gradient oracle:
  • $\mathbb{E}\left[G^{(i)}_{t}(\mathbf{x})\right]=\nabla f^{(i)}_{t}(\mathbf{x})$ for all $\mathbf{x}$;
  • $\mathbb{E}\left\|G^{(i)}_{t}(\mathbf{x})-\nabla f^{(i)}_{t}(\mathbf{x})\right\|^{2}\leq\sigma^{2}$ for all $\mathbf{x}$.
The next assumption connects the objective value with client representations. Client representations can take various forms, such as label distributions, input embeddings, or gradients.
@@
For a metric space $(X,\rho)$, let $r^{(i)}_{t}\in X$ be a representation of client $i$ at time $t$. Then there exists $\theta$ such that $\left|f^{(i)}_{t}(\mathbf{x})-f^{(j)}_{t}(\mathbf{x})\right|\leq\theta\cdot\rho\left(r^{(i)}_{t},r^{(j)}_{t}\right)$ for all $i,j,t$.
We next give the intuition of why the assumption is natural for the label-based and the embedding-based distnaces. If the concept $P(\text{output}|\text{input})$ doesn’t change between clients, every model achieves the same expected error on each class. For each class, assuming that the class data for different clients is sampled from the same distribution, we get the same expected loss on this class for different clients. Hence, the overall difference in the loss function is largely determined by the fraction of each class in the clients’ data. Similarly, if the client embedding faithfully captures information about its inputs, assuming the trained models depend on the inputs continuously, they have the same expected error on similar inputs.
@@
This assumption naturally holds in case when, for each client, only a small fraction of data points changes at every round. Finally, we assume that the clients are clusterable: there exist $K$ clusters so that within each cluster all points have similar representation.
@@
For completeness, we present the analysis of SGD convergence for functions satisfying $\mu$-PL condition.
Lemma 2.
@@
where in the last inequality we used $\eta\leq 1/L$. Using the fact that $f$ satisfies $\mu$-PL condition, we have $\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\geq 2\mu(f(\mathbf{x}_{t})-f(\mathbf{x}^{*}))$, giving
@@
The above result demosntrates how the objective improves with each round: namely, after $R$ iterations, the non-stochastic part of the loss improves by a factor which depends on $R$ exponentially. However, each data drift and each reclustering might potentially hurt the objective. Our next result bounds their effect on the loss.
Lemma 3.
@@
$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})=\left(f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})-f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})\right)+\left(f^{(i)}_{t+1}(\mathbf{c}_{k}^{*})-f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell}^{*})\right)$
@@
$f^{(i)}_{t+1}(\bar{\mathbf{c}}_{\ell})=f^{(i)}_{t+1}\Bigl{(}\operatorname*{avg}_{j\in\bar{C}_{\ell}}\mathbf{x}_{j}\Bigr{)}\leq\operatorname*{avg}_{j\in\bar{C}_{\ell}}f^{(i)}_{t+1}(\mathbf{x}_{j}),$ (1)
@@
where the last inequality holds since otherwise $f^{(i)}_{t}(\mathbf{c}_{k}^{*})>f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$, which is impossible since $\mathbf{c}_{k}^{*}$ is the minimizer of $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}$ and $\operatorname*{avg}_{j\in C_{k}}f^{(j)}_{t}(\mathbf{x}_{t}^{(i,*)})\leq f^{(i)}_{t}(\mathbf{x}_{t}^{(i,*)})+\theta\Delta$ by Assumptions D.
@@
Combining the above statements leads to our main convergence result.
Theorem 4.
@@
$\frac{1}{N}\sum_{k\in[K]}\sum_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t+1})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$ $\leq(1-\eta\mu)^{TR}\left(f(\mathbf{x}_{0})-f(\mathbf{x}^{*})\right)$
$\quad+\frac{L\eta}{2\mu}\left(\frac{\sigma^{2}+8L\theta\Delta}{M/K}+3\theta(\Delta+\delta)(1-\eta\mu)^{R}\right)$
@@
Next, by Lemma 3, the total objective increases by at most $3\theta(\Delta+\delta)$ after reclustering. Since during $R$ rounds the non-stochastic part of each objective decreases by a factor of $(1-\eta\mu)^{R}$, we have
@@
$\leq\frac{(1-\eta\mu)^{R}}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})+3\theta(\Delta+\delta)\right)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
$=(1-\eta\mu)^{R}3\theta(\Delta+\delta)+\sum_{i=0}^{R}\left(1-\eta\mu\right)^{i}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$
$\qquad+(1-\eta\mu)^{R}\frac{1}{N}\sum_{k\in[K]}|C_{k}|\operatorname*{avg}_{i\in C_{k}}\left(f^{(i)}_{t}(\mathbf{c}^{(k)}_{t})-f^{(i)}_{t}(\mathbf{c}^{(k,*)}_{t})\right)$
@@
By Lemma 2, the last (stochastic) term accumulates over all data drift events as $\frac{L\eta}{2\mu}\cdot\frac{\sigma^{2}+8L\theta\Delta}{M/K}$. By a similar reasoning, $3\theta(\Delta+\delta)$ accumulates as $\frac{L\eta}{2\mu}\cdot 3\theta(\Delta+\delta)(1-\eta\mu)^{R}$. And finally, $f(\mathbf{x}_{0})-f(\mathbf{x}^{*})$ term is multiplied by $(1-\eta\mu)^{R}$ at every data drift event, giving factor of $(1-\eta\mu)^{TR}$ over $T$ data drift events. ∎
Appendix B Additional background
In this section, we provide the background of clustered federated learning and quantify the impacts of data drift.
B.1 Clustered federated learning
FL deployment often involves hundreds to thousands of available clients participating in the training (Bonawitz et al., 2019). For instance, Huang et al. (2019) have developed mortality and stay time predictors for the eICU collaborative research database (Pollard et al., 2018) containing data of 208 hospitals and more than 200,000 patients. Although thousands of devices may be available in a given round, FL systems typically select only a subset for training. For example, Google sets 100 as the target number of clients per training round when improving keyboard search suggestions (Yang et al., 2018). This selection process ensures a reasonable training time and accounts for diminishing returns, where including more clients does not accelerate convergence (Xu et al., 2024).
Data heterogeneity is a major challenge in FL, as variations in local data volume and distribution among participants can lead to slow convergence and reduced model accuracy (Yang et al., 2021; Wen et al., 2023). Clustering is an effective strategy to mitigate the impact of heterogeneity by grouping statistically similar clients. Clients within the same cluster collaboratively train one cluster model, which then serves their inference requests. Prior works demonstrate that clustering achieve high accuracy and fast model convergence (Sattler et al., 2021; Mansour et al., 2020; Ghosh et al., 2020; Liu et al., 2023; Jothimurugesan et al., 2023).
Our experiments with the Fielding system prototype mainly use label distribution vectors as the client representation, as they handle both label and covariate shifts and are lightweight. Although previous studies have examined label distribution vectors as leverage to address heterogeneity (Zhang and Lv, 2021; Zhang et al., 2022; Lee and Seo, 2023; Lee et al., 2021), Fielding distinguishes itself from existing works. Federated learning via Logits Calibration (Zhang et al., 2022) and FedLC (Lee and Seo, 2023) are orthogonal and investigate building a better global model through client model aggregation and loss function improvements. FedLabCluster (Zhang and Lv, 2021) and Lee et al. (2021) do not explicitly discuss data drift handling. Moreover, we present a theoretical framework with a convergence guarantee under data drift and re-clustering.
B.2 Data drift affects static clustering
Drifts happen naturally when clients have access to non-stationary streaming data (e.g., cameras on vehicles and virtual keyboards on phones) (Koh et al., 2021; Bhardwaj et al., 2022; Chen et al., 2020; Nandi and Xhafa, 2022; Marfoq et al., 2023). For example, when training an FL model for keyboard suggestions, drifts occur when many clients simultaneously start searching for a recent event (widespread drift), a few clients pick up a niche hobby (concentrated intense drift), or a new term appears when a product is launched (new label added).
Data drift reduces the effectiveness of clustered FL as it increases data heterogeneity within clusters. We demonstrate this problem with the Functional Map of the World (FMoW) dataset (Christie et al., 2018). Recall from Section 3 that our preprocessed FmoW dataset contains time-stamped satellite images labeled taken on or after January 1, 2015, and each unique UTM zone metadata value corresponds to one client (302 clients in total). A day’s worth of data becomes available every two training rounds, and clients maintain images they received over the last 100 rounds.
As mentioned in Section 2, we use mean client distance as the intra-cluster heterogeneity. Note that instead of first averaging within each cluster and then finding the mean of those cluster averages, we use the overall mean across all clients. This helps us avoid biased results arising from imbalanced cluster sizes. Note that in the baseline case without any clustering, we consider all clients as members of a "global cluster".
We cluster clients based on their label distribution vectors at the starting round and track the overall heterogeneity by measuring the mean client distance described above. Figure 1 shows that initially (e.g., round 1), clustering reduces the per-cluster heterogeneity compared with no clustering (i.e., putting all clients in a global set). However, as data distribution shifts over time, static clustering increases per-cluster heterogeneity and soon gets close to the no clustering case at round 322.
Appendix C Implementation
Figure 11: The system architecture and workflow of Fielding. Blue arrows represent communication between clients and the coordinator. Black arrows indicate the workflow among different components on the same machine.
We implemented Fielding in Python on top of FedScale (Lai et al., 2022), a state-of-the-art open-source FL engine. We follow FedScale’s design of having one centralized coordinator and client-specific executors. Figure 11 illustrates the end-to-end workflow of the Fielding system prototype with label distribution vectors as client representations. At the beginning of each training round, 1 client executors register with the coordinator to participate and report their local data distribution optionally. The coordinator’s cluster manager maintains clients’ distribution records and 2 moves clients to the closest cluster when their reported distribution drifts. After moving drifted clients individually, the cluster manager measures center shift distances and either triggers global re-clustering when any cluster center shifts significantly or finalizes clusters membership. In the case of global clustering, we use $K$-means clustering and determine $K$ using the silhouette method (Rousseeuw, 1987). The initial model of each newly created cluster is set as the average of its clients’ previous cluster model (see Algorithm 2).
Appendix D Experimental settings